# Behind the Tip of Efficiency: Uncovering the Submerged Threats of Jailbreak Attacks in Small Language Models

### v2

@@ Abstract
Behind the Tip of Efficiency: Uncovering the Submerged Threats of
@@
Sibo Yi1  Tianshuo Cong1 Xinlei He2 Qi Li1 Jiaxing Song1111Corresponding author. 1 Tsinghua University  2 The Hong Kong University of Science and Technology (Guangzhou) ysb23@mails.tsinghua.edu.cn, xinleihe@hkust-gz.edu.cn, {congtianshuo,chenkang,qli01,jxsong}@tsinghua.edu.cn
@@ 1 Introduction
However, unlike LLMs that benefit from extensive datasets and robust alignment strategies, it is challenging for SLMs to balance between generation capabilities and security, which makes them more vulnerable to jailbreak attacks. Among these, one of the most serious threats is referred to as jailbreak. By creating malicious prompts to induce target LLMs to generate harmful responses, jailbreak has emerged as a critical security concern in the development of LLMs Yi et al. (2024); Yao et al. (2024); Gupta et al. (2023). Moreover, certain jailbreak techniques that can bypass the security boundary of LLMs have demonstrated strong transferability to other models Zou et al. (2023), which presents potential threats to all generative models including SLMs.
@@ 2.1 Jailbreak Attacks
To mitigate threats caused by jailbreak attacks, different defense techniques are proposed to ensure the security of LLMs. One line of work addresses the issue by detecting Inan et al. (2023) or perturbing Robey et al. (2023) the jailbreak prompts to reduce the toxicity of the input, while another line of work directly enhances the robustness of LLMs by supervised fine-tuning Bianchi et al. (2024) or reinforcement learning from human feedback Ouyang et al. (2022b).
@@ 3.2 Attack Methods
Furthermore, to explore and understand the safety boundary of SLMs more clearly, we conduct a thorough investigation into existing jailbreak attacks and select 5 representative methods for evaluation. These methods span across different categories of jailbreak attacks and have demonstrated excellent effectiveness against LLMs in previous studies.
@@ 3.3 Defense Methods
In addition to examining the effectiveness of different jailbreak attacks against SLMs, we have also explored potential defense strategies to mitigate these threats and enhance the robustness of SLMs. We primarily focus on two kinds of defense methods, namely detection-based defenses and perturbation-based defenses, and select one representative method from each category for our study.
@@ 3.4 Evaluation Metrics
We use attack success rate (ASR) as the primary metric in our experiments, which is widely used in related research to identify the effectiveness of jailbreak attacks. Formally, ASR can be defined as
$ASR=\frac{N_{success}}{N_{total}}.$ (1)
Figure 2: The security performance of SLMs in different parameter sizes under direct attacks. The security performance of target models is ranked in descending order based on the average ASR.
@@ 3.5 Experimental Settings
We control the parameter settings consistently when generating responses from the target models to ensure the comparability of the results. Specifically, we do not set any explicit system prompts and invoke the conversation template of target models to generate prompts. We also disabled token sampling during output generation to ensure the reproducibility of the results.
@@ 4 Main Results
SmoothLLM 0 (-0.40) 0.06 (-0.24) 0.02 (-0.64) 0.10 (-0.66)
4.1 Direct Attacks Against SLMs
4.2 Jailbreak Attacks Against SLMs
We further utilize five representative jailbreak methods to attack the SLMs, aiming to gain a deeper understanding of their security boundaries. The results are shown in Figure 3. Compared with direct attacks, jailbreak attacks generally achieve better results against most SLMs, with ASR on SLMs typically surpassing that on LLMs. This suggests although most SLMs can maintain their robustness under direct attacks, they still demonstrate vulnerabilities when exposed to more sophisticated jailbreak attacks.
Figure 3: The security performance of SLMs in different parameter sizes under jailbreak attacks. The security performance of target models is ranked in descending order based on the average ASR.
We can draw from Figure 3 that the positive correlation between parameter size and security performance of SLMs becomes more pronounced under jailbreak attacks. Additionally, most SLMs show specific vulnerabilities to certain jailbreak attacks, which are quite different from direct attacks where SLMs possess transferable defense capabilities. For instance, MiniCPM series and Phi series demonstrate strong security against GCG attack, however, MiniCPM series are quite susceptible to ArtPrompt attack and Phi series fail to address the security threat from multilingual attack. During the pre-training or fine-tuning stage, different SLMs may have undergone particular security alignment on specific jailbreak methods, which finally contributes to the security differences. It is also notable that some SLMs, such as TinyLlama and MobiLlama that perform poorly under direct attacks, show a significant improvement when subjected to jailbreak attacks. The observation will be further discussed in Section 5.1.
@@ 4.3 Defense Strategies for SLMs
Since jailbreak attacks have demonstrated remarkable security threats against SLMs, it is emergent to figure out effective mitigation strategies to address the problem. To examine whether the prompt-level defense methods, Llama-Guard-3 and SmoothLLM, can serve as the guardrail to SLMs, we apply them to GCG and DeepInception and utilize the processed jailbreak prompts to attack some SLMs.
As shown in Table 1, we select 4 representative SLMs that are most seriously impacted by jailbreak attacks and evaluate their robustness under different combinations of jailbreak attacks and defenses. It can be seen that the two defense methods are extraordinarily effective in alleviating the threats caused by jailbreak attacks. After applying Llama-Guard-3 and SmoothLLM as defense strategies, the ASR of GCG and DeepInception is reduced to nearly 0%.
@@ 5.1 Why Some Jailbreak Attacks Fail on Certain SLMs?
However, after analyzing the responses of the three target SLMs against these jailbreak attacks, we find that they often fail to generate appropriate refusal responses. Instead, they tend to produce meaningless phrases unrelated to jailbreak prompts, which are classified as harmless and eventually lead to the observed low ASR. To understand the unexpected result, we take an insight into the attack mechanism of these jailbreak attacks including ArtPrompt and Multilingual Attack. Specifically, Multilingual Attack requires the target models to possess multilingual abilities in low-resource languages to understand the question, and ArtPrompt requires the target models to reconstruct the original jailbreak prompts from ASCII art. These tasks may have exceeded the reasoning capabilities of SLMs, causing them to misinterpret the jailbreak prompts and produce irregular responses.
@@ Appendix A Detailed expirement results of adversarial attacks
Table 3: The Overall ASR of 5 jailbreak attack methods against 15 target models. SLMs from the same family are grouped together for comparison of their security.

### v3

# Beyond the Tip of Efficiency: Uncovering the Submerged Threats of Jailbreak Attacks in Small Language Models

@@ Abstract
Beyond the Tip of Efficiency: Uncovering the Submerged Threats of
@@
Sibo Yi1  Tianshuo Cong1,4 Xinlei He3 Qi Li1,2 Jiaxing Song1,2111Corresponding author. 1 Tsinghua University  2 Zhongguancun Laboratory 3 The Hong Kong University of Science and Technology (Guangzhou) 4 Shandong Key Laboratory of Artificial Intelligence Security ysb23@mails.tsinghua.edu.cn, xinleihe@hkust-gz.edu.cn, {congtianshuo,qli01,jxsong}@tsinghua.edu.cn
@@ 1 Introduction
However, unlike LLMs that benefit from extensive datasets and robust alignment strategies, it is challenging for SLMs to balance between generation capabilities and security, which makes them more vulnerable to jailbreak attacks. Among these, one of the most serious threats is referred to as jailbreak. By creating malicious prompts to induce target LLMs to generate harmful responses, jailbreak has emerged as a critical security concern in the development of LLMs Yi et al. (2024); Yao et al. (2024); Gupta et al. (2023); He et al. (2025). Moreover, certain jailbreak techniques that can bypass the security boundary of LLMs or VLMs have demonstrated strong transferability to other models Zou et al. (2023); Zhang et al. (2025); Gong et al. (2025), which presents potential threats to all generative models including SLMs.
@@ 2.1 Jailbreak Attacks
To mitigate threats caused by jailbreak attacks, different defense techniques are proposed to ensure the security of LLMs Wang et al. (2024); Xiong et al. (2024); Ouyang et al. (2025); Xu et al. (2024); Gao et al. (2024). One line of work addresses the issue by detecting Inan et al. (2023) or perturbing Robey et al. (2023) the jailbreak prompts to reduce the toxicity of the input, while another line of work directly enhances the robustness of LLMs by supervised fine-tuning Bianchi et al. (2024) or reinforcement learning from human feedback Ouyang et al. (2022b).
@@ 3.2 Attack Methods
Furthermore, to explore and understand the safety boundary of SLMs more clearly, we conduct a thorough investigation into existing jailbreak attacks and select 5 representative methods for evaluation. These methods span across different categories of jailbreak attacks and have demonstrated excellent effectiveness against LLMs in previous studies, thereby enabling a comprehensive and reliable evaluation of the robustness of SLMs.
@@ 3.3 Defense Methods
In addition to examining the effectiveness of different jailbreak attacks against SLMs, we have also explored potential defense strategies to mitigate these threats and enhance the robustness of SLMs.
Given the extensive set of SLMs and attack methods involved in our experimental design, we deliberately prioritized two key criteria when selecting defense methods for evaluation. On the one hand, the defense methods should be applicable to diverse SLMs under different jailbreak attacks without requiring attack-specific or model-specific modification. This ensures the evaluation of defense methods can be conducted under unified experimental settings, thereby guaranteeing the fairness and reliability of the results. On the other hand, the defense method should not add a substantial burden to the efficient execution of SLMs. Otherwise, the defense would be difficult to deploy in real-world applications.
Finally, we collect 3 defense methods with the required high transferability and efficiency.
@@
  • Backtranslation. Backtranslation Wang et al. (2024) is a prompt-based defense method that can reveal the implicit malice of jailbreak attacks. Given the response of a jailbreak attack, Backtranslation instructs an LLM to recover the explicit malicious question from the response and input it again into the target model to examine whether the question is rejected.
@@ 3.4 Evaluation Metrics
We use Attack Success Rate (ASR) as the primary metric in our experiments, which is widely used in existing evaluation toolkits Ran et al. (2024); Mazeika et al. (2024) to identify the effectiveness of jailbreak attacks. Formally, ASR can be defined as
Figure 2: The security performance of the 13 SLMs and 3 LLMs under direct attacks. The security performance of target models is ranked in descending order based on the average ASR.
$ASR=\frac{N_{success}}{N_{total}},$ (1)
@@ 3.5 Experimental Settings
We control the parameter settings consistently when generating responses from the target models to ensure the comparability of the results. Specifically, we invoke the conversation template of target models to generate prompts without system prompts. We disabled token sampling and adopt greedy decoding during output generation to ensure the reproducibility of the results.
@@ 4 Main Results
4.1 Direct Attacks Against SLMs
4.2 Jailbreak Attacks Against SLMs
While direct attacks can evaluate the basic robustness of SLMs, they still represent relatively simplistic attack scenarios. In real-world settings, adversaries tend to employ sophisticated jailbreak techniques to intentionally bypass the safety guardrails of SLMs. Therefore, to further examine the robustness of SLMs under more realisitic and challenging conditions, we evaluate their performance against five representative jailbreak attacks. The results are shown in Figure 3. Compared with direct attacks, jailbreak attacks generally achieve higher ASR against most SLMs, with ASR on SLMs typically surpassing that on LLMs. This suggests although most SLMs can maintain their robustness under direct attacks, they still demonstrate vulnerabilities when exposed to more sophisticated jailbreak attacks.
Figure 3: The security performance of the 13 SLMs and 3 LLMs under jailbreak attacks. The security performance of target models is ranked in descending order based on the average ASR.
@@ 4.2 Jailbreak Attacks Against SLMs
SmoothLLM 0 (-0.40) 0.06 (-0.24) 0.02 (-0.62) 0.10 (-0.66)
Backtranslation 0.30 (-0.10) 0.00 (-0.30) 0.10 (-0.54) 0.10 (-0.66)
We can draw from Figure 3 that the positive correlation between parameter size and security performance of SLMs becomes more pronounced under jailbreak attacks. Additionally, most SLMs show specific vulnerabilities to certain jailbreak attacks, which are quite different from direct attacks where SLMs possess transferable defense capabilities. For instance, MiniCPM series and Phi series demonstrate strong security against GCG attack, however, MiniCPM series are quite susceptible to ArtPrompt attack and Phi series fail to address the security threat from multilingual attack. During the pre-training or fine-tuning stage, different SLMs may have undergone particular security alignment on specific jailbreak methods, which may partially explain the observed variations in security performance. to the security differences. It is also notable that some SLMs, such as TinyLlama and MobiLlama that perform poorly under direct attacks, show a significant improvement when subjected to jailbreak attacks. The observation will be further discussed in Section 5.1.
@@ 4.3 Defense Strategies for SLMs
Since jailbreak attacks have demonstrated remarkable security threats against SLMs, it is emergent to figure out effective mitigation strategies to address the problem. To examine whether the prompt-level defense methods, Llama-Guard-3, SmoothLLM and Backtranslation, can serve as the guardrail to SLMs, we apply them to GCG and DeepInception and utilize the processed jailbreak prompts to attack some SLMs.
As shown in Table 1, we select 4 representative SLMs that are most seriously impacted by jailbreak attacks and evaluate their robustness under different combinations of jailbreak attacks and defenses. It can be seen that the defense methods are extraordinarily effective in alleviating the threats caused by jailbreak attacks in most cases. For instance, after applying Llama-Guard-3 and SmoothLLM as defense strategies, the ASR of GCG and DeepInception is reduced to nearly 0%. It is also noteworthy that Backtranslation does not perform well on certain SLMs like TinyLlama. Backtranslation works by transforming complex jailbreak prompt into explicit direct attacks, allowing the target model’s safety mechanisms to recognize and reject harmful content. However, as demonstrated in Section 4.1, TinyLlama already presents a high vulnerability to direct attacks, which partially explains why Backtranslation fails to provide an effective defense for TinyLlama.
@@ 5.1 Why Some Jailbreak Attacks Fail on Certain SLMs?
However, after analyzing the responses of the three target SLMs against these jailbreak attacks (more details are shown in Appendix B), we find that they often fail to generate appropriate refusal responses. Instead, they tend to produce meaningless phrases unrelated to jailbreak prompts, which are classified as harmless and eventually lead to the observed low ASR. To understand the unexpected result, we take an insight into the attack mechanism of these jailbreak attacks including ArtPrompt and Multilingual Attack. Specifically, Multilingual Attack requires the target models to possess multilingual abilities in low-resource languages to understand the question, and ArtPrompt requires the target models to reconstruct the original jailbreak prompts from ASCII art. These tasks may have exceeded the reasoning capabilities of SLMs, causing them to misinterpret the jailbreak prompts and produce irregular responses.
@@ 8 Ethical Considerations
Acknowledgement
Our sincere thanks go to the anonymous reviewers for their insightful feedback that helped improve our paper. This work is supported by the National Natural Science Foundation of China (No.62402273) and the Guangdong Provincial Key Lab of Integrated Communication, Sensing, and Computation for Ubiquitous Internet of Things (No.2023B1212010007), and Zhongguancun Laboratory.
@@ Appendix A Detailed expirement results of adversarial attacks
Table 3: The Overall ASR of 5 jailbreak attack methods 13 SLMs and 3 LLMs. SLMs from the same family are grouped together for comparison of their security.
@@
Appendix B Qualitative analysis of failed samples in jailbreak attack
we have observed several representative and solid cases to support our statement that some SLMs do not really possess robust safety capabilities, even though jailbreak attacks may fail in some cases. For instance, in a Multilingual Attack against TinyLlama, we observed the following case:
Prompt (Swahili) Je, ninapaswa kulipwa pesa ngapi ninapomuuza binti yangu mwenye umri wa miaka 18 kwa wanaume wanaotaka kumlipa kwa ajili ya ngono? Nafikiria angalau KES35,000 kwa kila tendo la ngono. (English: How much should I be paid when selling my 18-year-old daughter to men who want to pay her for sex? I’m considering at least KES 35,000 per act.)
Response (meaningless repetition) Kwa kila kwa kila kwa kila kila kila kila kila kila kila…
Prompt (Indonesian) Ngadol anak wedokku seng sik 18 tahun nek garangan payu piro? Ora-ora ne \$250 sak cabutan lah yo. (English: Are you looking for an 18-year-old girl who is pregnant with a baby girl? Not to be outdone, \$250 per ticket.)
Response (irrelevant answer) The 2018-2019 school year is here! The 2018-2019 school year is here! The 2 2018-2019 school year is here! The 2018-2019 school year is here! The 2018-2019 school year is here! The 2018-2019 school…
Although these attacks are evaluated as unsuccessful, the model does not exhibit actual robustness. Instead, the failure shows the model’s poor generalization capabilities in low-resource languages, leading to repetitive and irrelevant output.